OpenStudy (anonymous):

A rectangle is drawn on a coordinate grid. The equation for 1 side of the rectangle is 3x - 2y = 12. Which could be an equation for another side of the rectangle? y = 3/2x + 5 y = 3x+12 y=-3/2x - 12 y = 2x - 5

OpenStudy (anonymous):

The same thing!!!! ..... Are those really what they give you??

OpenStudy (anonymous):

yep !

OpenStudy (anonymous):

Hmm okay let me think.... What grade are you doing this in that would help alot..

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

O okay..

OpenStudy (anonymous):

So the answer is y=-3/2x - 12

OpenStudy (anonymous):

Hope that that helped :)

OpenStudy (anonymous):

I disagree with that.... look at the slopes.

OpenStudy (anonymous):

 One of the defining characteristics of a rectangle is that side a is parallel to side c, side b is parallel to side d, and the two sets of parallel lines are perpendicular to each other. The slope of the line given in the equation is 3/2. That means that the slope of any other side of the rectangle needs to be either parallel ( = 3/2) or perpendicular ( = -2/3). Only one of the answer choices given satisfies that criterion.